OpenStudy (anonymous):

simplify & write in lowest terms: 6xy/8xky divided by 3z/8k

OpenStudy (anonymous):

\[\frac{ 6xy }{ 8xky } \div \frac{ 3z }{ 8k }\]

OpenStudy (anonymous):

\[ \frac{ 6xy }{ 8xky } \div \frac{ 3z }{ 8k } = \frac{ 6xy }{ 8xky } \times \frac{ 8k }{ 3z } \]Can you multiply fractions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

6xy/8xky / 3z/8k first deal with the numbers 3xy/8xky / 3z/4k then deal with the letters xy and y cancel in the first part so you have 3/8k / 3z/8k then you cross multiply by the reciprocal and your answer is z

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):